OpenStudy (anonymous):

An 80.0-gram sample of a gas was heated from 25 °C to 225 °C. During this process 346 J of work was done by the system (which was stationary), and its internal energy increased by 7765 J. What is the specific heat capacity of the gas? i just want know whether work is being done. I understand that no work is done on an object if that object remains stationary. in this case the stationary system is doing the work. Is it though?

OpenStudy (kainui):

Work is being done. I'm not entirely sure what they mean by stationary, since there has to be some form of expansion work going on here when you increase the temperature of a gas.

OpenStudy (ghazi):

@hammed when we talk about expansion of gases and thermodynamics we don't see displacement of object and we don't apply w= force * distance now here you have to use Q=U+W where Q is heat in the process, U is internal energy and W is work also Q= m*C* change in temperature

OpenStudy (ghazi):

\[Q= \Delta U+ W\] and \[Q= m*C * \Delta T\] C is specific heat capacity

OpenStudy (ghazi):

\[m*C* \Delta T= 7765+ 346\] and \[\Delta T= (225-25)\]

OpenStudy (anonymous):

thanks